# Advanced Architectures Integrated with Agentic AI for Next-Generation Wireless Networks

### v2

@@ IV Conclusion
Additionally, we explored the Neural Radio Protocol Stacks using Agentic AI that highlights the potential for customizable and efficient radio protocols in future networks. We conducted an experiment for parameter optimizing in the context of V2X with and without AI agents. We designed 3 AI agents to provide optimized parameters recursively based on the current state and evaluated the performance using PDR, link robustness, retransmission rate, adaptability, and latency. We show that the parameter optimization with Agentic AI is far better than the conventional techniques. Together, these advancements represent critical steps toward realizing the 6G vision, where networks are more adaptable, intelligent, and capable of supporting diverse applications with greater efficiency and flexibility.

### v3

@@ V Conclusion
Additionally, we explored the Neural Radio Protocol Stacks using Agentic AI that highlights the potential for customizable and efficient radio protocols in future networks. We conducted an experiment for parameter optimizing in the context of V2X with and without AI agents. We designed 3 AI agents to provide optimized parameters recursively based on the current state and evaluated the performance using PDR, link robustness, retransmission rate, adaptability, and latency. We show that the parameter optimization with Agentic AI is far better than the conventional techniques.
As 6G standardization activities gain momentum in organizations such as 3GPP, ITU-R, and IEEE, the integration of Agentic AI principles into future specifications will be crucial. Our work provides technical foundations that can inform these standardization efforts, particularly in the areas of network architecture, constrained AI operations, and dynamic resource optimization. The experimental validation demonstrates the tangible benefits of Agentic AI approaches, supporting their consideration in upcoming 6G standards.